OpenStudy (anonymous):

A marksman hits a target with probability 5/6 Assuming independence for successive firings, find the probabilities of getting (a) One miss followed by two hits. (b) Two misses and one hit (in any order).

OpenStudy (zarkon):

a)\[\frac{1}{6}\cdot\frac{5}{6}\cdot\frac{5}{6}\]

OpenStudy (zarkon):

\[_3C_1\left(\frac{5}{6}\right)^1\left(\frac{1}{6}\right)^2\]

OpenStudy (zarkon):

for b

OpenStudy (anonymous):

so i just multiply 5/6 and (1/6)^2

OpenStudy (zarkon):

3*(5/6)*(1/6)^2

OpenStudy (gw2011):

A) (1/6)(5/6)(5/6)=25/216 B) (1/6)(1/6)(5/6)=5/216

OpenStudy (zarkon):

(b) is 5/72